OpenStudy (anonymous):

how many 5-person committees are possible from a group of 9 people if: a. there are no restrictions? b. both Jim and Mary must be on the committee? c. Either Jim or Mary (but not both) must be on the committee?

OpenStudy (anonymous):

45

OpenStudy (anonymous):

first one is \(\binom{9}{5}=\frac{9\times 8\times 7\times 6}{4\times 3\times 2}=156\)

OpenStudy (anonymous):

Second one is \(\binom 9 3 = 84 \)

OpenStudy (anonymous):

Third one is \( 2 \times \binom 8 4 =140\)

OpenStudy (phi):

I would amend Fool's answer for (b). If Jim and Mary must be on the committee, then you only have 7 other people from which to choose the remaining 3. so \[\left(\begin{matrix}7 \\ 3\end{matrix}\right)=35\]

OpenStudy (anonymous):

OOoooooooooooooppos!! Yes, They are already selected so no need to count them in the total number.